OpenStudy (anonymous):

If you have 200 grams of a radioactive sample, how many grams would be left after three half-lives? Select one: a. 200 b. 150 c. 50 d. 25

OpenStudy (anonymous):

200 * 1/2 = 100 100 *1/2 = 50 50 * 1/2 = 25 D. 25

OpenStudy (anonymous):

thanks wish i can give a medal

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i got a lot if you dont mind

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me post them

OpenStudy (anonymous):

Radioactive decay can be used to measure: Select one: a. time b. length c. circumference d. density

OpenStudy (anonymous):

A graph of exponential decay has ____________ on the x axis and ____________ on the y axis. Select one: a. Time; number of radioactive atoms b. Radioactivity; time c. Electrons; time d. Time; neutrons

OpenStudy (anonymous):

A, and i don't know about that 2nd one

OpenStudy (anonymous):

thank you